# Boosting Mixed-Initiative Co-Creativity in Game Design: A Tutorial

### v1

@@ 3.2.2. RL Brush
Delarosa et al. (Delarosa et al., 2021) presented a mixed-initiative tool for designing tile-based levels, with a version specifically tailored for the puzzle game Sokoban. RL Brush (Fig. 4) uses a set of AI models to interactively suggest incremental modifications while the human designer is editing the level. The system presents four suggestions at a time generated by reinforcement-learning-based agents: a narrow agent that can change a specified tile at each step, a turtle agent that can move to a next tile in the map grid and modify passed tiles, a wide agent that has control to edit any tile in the map, and a majority agent that suggests tile mutations if the majority of the other three agents have suggested that same mutation. The display of new suggestions is triggered by user interactions only. The user can accept any of these suggestions or override them by manually editing the grid for the current level. The tool also allows the user to fine-tune the performance of the AI models by providing two hyperparameters for the user to control. The step parameter defines the number of interactions for the manager of the models to recursively call itself so that a higher step value allows for more than one modification to the map. The tool radius parameter controls how big is the window of tiles that the artificial agents can view and modify. By generating modifications one step at a time, this approach results in a creative process in which participants gradually build the level tile by tile.
Figure 4. The RL Brush tool (Delarosa et al., 2021). Top - the step and tool radius parameters; left - the toolbar to edit, erase, delete, play, and save the current level; middle - the editable grid of the current level and the four AI suggestions on top; right - the tileset. (Available at https://rlbrush.app/)
@@
Figure 5. The miWFC tool (Langendam and Bidarra, 2022). Left - a list with several example input images with the selected one displayed; middle - the panel for manipulating tile properties; right - the current image accompanied by a timeline indicating the progress of the generation and a number of controls for adding markers and navigating through the timeline. The screen for direct manual editing is accessible from the brush button in the bottom right corner. (Available at https://github.com/ThijmenL98/miWFC)
@@ 3.3.1. Pitako
Machado et al. (Machado et al., 2019a, b) developed a tool for designing tile-based games that applies the concept of a recommender system to suggest mechanics and dynamics learned from existing games. The system is built on Video Game Description Language (VGDL), which considers sprites as objects with a graphical representation and an associated behavior. The creative process can begin by adding sprites to the game in development. For instance, the user can define a sprite behavior as a ShootAvatar and then attach the image of the hero to it. Having a catalog of a vast number of games, the system can provide suggestions of sprites to the designer, who is free to accept them or not (Fig. 6). If the user accepts a recommended sprite, it is imported to the game sprite set along with any others it may carry (e.g. picking a sprite that can shoot will also import the sprite for the projectile behavior). In addition, the system also suggests interactions for pairs of sprites (e.g. enemy sprite kills avatar sprite), taking its cues from the existing games in the catalog. If the user selects a recommendation, the interaction is automatically implemented in the user’s game.
This mixed-initiative approach is primarily targets content creation within the class of game behaviors. However, it also extends to the game space and game design classes by enabling the placement of sprites on the level grid and the definition of game termination rules. Still, it only serves as a support tool for these two classes as the computational agent lacks initiative.
Figure 6. Sprite Recommendation in the Pitako tool (Machado et al., 2019a). Top - a menu featuring tabs for the several tasks within the game design process; left - the game’s sprite set, initially containing only a default background sprite; center - the sprite property editor, enabling the selection of an image and the configuration of its behavior; right - a list of the system-generated suggestions. (Available at https://github.com/tiago-lam/Recommender-System-for-Game-Design)
@@ 3.4.1. EvolvingBehavior
Partlan et al. (Partlan et al., 2022) introduced a tool for crafting non-player character (NPC) behavior by evolving behavior trees in Unreal Engine 4. The process begins with the human designer providing an initial draft of a tree and other materials, such as a library of pre-defined nodes and templates for generating additional nodes. The system can then modify and generate behavior trees, even if the initial tree is empty or incomplete. In order for the evolution of behavior trees to meet the goals of the human designer, they must define a fitness function by setting fitness measurement keys. The user also needs to define several parameters of the evolutionary process. Given that the evolved trees can be manually edited, in practice, the human designer can modify them and use them as input for a new round of evolution, engaging in an iterative process. However, this work was not selected for further analysis in section 4. This is because it is currently in a state where user interaction with the tool is too complex for a regular designer without expert knowledge of the system and coding skills to easily work with it. Particularly, it is challenging to define the fitness function and adjust all the parameters without prior knowledge. Additionally, the generation and subsequent simulation of behavior trees are time-consuming. The delay would likely discourage designers from participating in a new evolutionary process.
@@ 3.5.1. Story Designer
The Evolutionary Dungeon Designer (EDD) is a mixed-initiative tool for designing 2D dungeons and the rooms that compose them. EDD is part of an ongoing research project and has been progressively expanded over several works (Baldwin et al., 2017a, b; Alvarez et al., 2018a, b, 2019, 2021). As a next step, Alvarez et al. (Alvarez et al., 2022) built an approach to design narrative structures integrated in EDD. Story Designer (Fig. 7) uses tropes in the form of nodes that can be interconnected in a graph to construct a complete narrative structure. The creative process begins with the system providing a basic sample graph HERO → CONFLICT → ENEMY. Users can manually edit the narrative graph by adding and connecting nodes representing characters, their roles, goals, and events. This triggers an evolutionary algorithm that generates alternative suggestions of narrative structure in real time that the user can inspect and choose to replace the current narrative design. The algorithm is an adaptation of MAP-Elites, which evaluates the set of solutions according to metrics such as interestingness and coherence. The system displays the suggestions in a matrix arranged by their values across specified dimensions of interest. Solutions with optimal fitness are shown in dark green, while dark red cells mean the best solution for those values is still not feasible. The human designer can visualize two dimensions at a time a switch between them as needed. As Story Designer is linked to the level design tool of EDD, the suggested narrative graphs are constrained by the existing content in the level’s map.
Figure 7. The Story Designer tool (Alvarez et al., 2022). Center - The editor for the main narrative graph; right - the suggestion grid, the list of dimensions that the user can choose for the grid, the suggestion they are hovering, and the selected suggestion; bottom - display of fitness metrics of the main graph and of the suggestions. (Available at https://github.com/mau-games/eddy)
@@
Figure 8. The 1001 Nights game (Sun et al., 2022). Left - a dialogue interface containing text blocks from the story told by the player and the King, along with a text box at the bottom for the player to write; right - a screen featuring the game characters, in which the player can collect the generated items and add them to the inventory at the bottom; bottom right corner: the button to initiate the battle. (Available at https://cheesetalk.itch.io/one-thousand-and-one-night)
@@ 3.6. Game Design
There are few MI-CCy approaches that address high-level concepts, such as the rules and objectives of a game. Most existing tools are already tailored for a specific game or game genre. So far, when the creative task of involves ideating a novel game, it is still uncommon to grant decision-making authority to the computational agent, thus leaving it with a passive role in the process. Nevertheless, there are some works that deviate from this pattern. Among the works we have reviewed, CADI and Gamika fall into this category but were not selected. CADI (Mobramaein et al., 2018) is a tool that employs a conversational interface for agents to design variations of the game Pong. Gamika (Nelson et al., 2017) is a system for creating 2D physics-based games for mobile devices, which parametrizes the design space into 284 features that the user can manipulate.
@@ 3.6.1. Germinate
For instance, a user interested in creating a stress-themed game might define entities such as “Friend” and “Work”, a resource called “Stress”, and the relationships and triggers they want to showcase. This example is illustrated in Fig. 9. The user can explore the several generated games to assess if there are any interesting results and then make adjustments to the properties to better reflect their intent, allowing for iterative refinement of the game design.
This approach is focused on starting from high-level rhetorical goals and then working on details to meet those goals. To articulate the conceptual essence of the game, Germinate also relies on the creation of a game system and game behaviors.
Figure 9. The Germinate tool (Kreminski et al., 2020). Left - properties specified by the user; center - a playable version of one of the generated games (a single game is shown at a time); right - the properties of the generated game. (Available at https://github.com/ExpressiveIntelligence/Germinate)
@@ 4. Mixed-initiative Classification
Deterding et al. (Deterding et al., 2017) emphasize how the human and the computer iteratively respond to each other through a variety of possible actions. Following this idea, Spoto and Oleynik (Spoto and Oleynik, 2017) introduced a framework that allows the comparison of mixed-initiative systems regarding the flow of the creative process. The flow is visualized through a graph that considers seven subprocesses of the creative process, namely ideate, constrain, produce, suggest, select, assess, and adapt, which can be executed by both the human and the computer. The graph illustrates the sequence of actions through directional links, usually outlining internal interactive loops. Muller et al. (Muller et al., 2020) expand on this framework, addressing mixed-initiative scenarios with generative AI systems. Grabe et al. (Grabe et al., 2022) also adapt the framework to co-creativity with Generative Adversarial Networks (GANs), identifying four primary patterns of interaction between humans and GANs. Focused on interaction design in co-creative systems, Rezwana and Maher (Rezwana and Maher, 2021, 2023) presented COFI, a framework that describes several interaction components to represent the space of possibilities for interaction between humans and AI and with the shared creative content. We consider these frameworks to be valuable, making it possible to distinguish between approaches in terms of how the creative process admits initiative from both agents involved and how they interact with each other. However, they make little to no mention of the degree of mixed-initiative, the real impact each agent has on the creative process and how much their collaboration is facilitated.
@@ 4.1. MI-CCy Quantifier: A Framework for Mixed-initiative Co-creativity Quantification
Liapis et al. (Liapis et al., 2016) consider it ingenuous to expect human and computational agents to be equal partners in the creative process. We admit that participants involved in co-creation can never be exactly equal throughout the various stages of the process, which is also true for human-human co-creativity, as each person has their own personality, knowledge, capabilities, and past experience. Nonetheless, we argue that such should be a goal for future mixed-initiative tools. Liapis et al. (Liapis et al., 2016) reiterate that current approaches favor human control, with the user having the final say on the decisions taken. We hope that new approaches can challenge this view. It is our belief that working as colleagues in a process in which collaboration is continuous and facilitated allows the human and the machine to learn from each other productively and can maximize the capabilities of both. With this perspective in mind, we present the MI-CCy Quantifier, a framework for comparing systems with respect to the degree of mixed-initiative co-creativity. The classification is made independently of the process flow. In this way, the analyzed parameters are generalized and applicable to any system, even if the processes are very different in their form. This model allows for a more direct visualization of which agent has a greater impact and how seamless their collaboration is. Figure 10 shows the framework representing a scenario with a very high degree of mixed-initiative co-creativity.
@@
We introduce four parameters that allow us to classify the influence that the human and the computer have on the creative process: Initial Setting, Initiative, Evaluation, and Final Decision. Each of these parameters is assessed along a spectrum, with the human at one end and the computer at the other. We consider that the more equal (towards the middle of the spectrum), the higher the degree of mixed-initiative. The framework has three more parameters that classify the approach on how much it promotes seamless collaboration. They are Task Assignment, Intervention Pace, and Explainability. Each of these parameters comprises its own scale, which is displayed vertically. We consider that the higher up the scale, the more seamless the collaboration between the human and the machine will be and, consequently, the greater the degree of co-creativity.
@@
It should be noted, however, that just because one system has a high degree of mixed-initiative co-creativity over another does not mean that the approach is better than the other. It only tells us that the human agent and the computational agent are more equal partners and are equipped for a high level of collaboration. Such may or may not be desirable, depending on the intended applicability of the system. If the user already has an established idea and only wants help to execute it, they do not expect to have a system that wants to have a strong initiative in the conceptual phase of the process. Nevertheless, in this tutorial, we encourage researchers interested in developing new mixed-initiative co-creative systems to seek to implement an approach with a higher degree of MI-CCy since there is still little exploration in the current state of the art of the potential of the computer as a “true” colleague in the creative process.

### v2

@@ 3.2.2. RL Brush
Delarosa et al. (Delarosa et al., 2021) presented a mixed-initiative tool for designing tile-based levels, with a version specifically tailored for the puzzle game Sokoban. RL Brush (Fig. 4) uses a set of AI models to interactively suggest incremental modifications while the human designer is editing the level. The system presents four suggestions at a time generated by reinforcement-learning-based agents: a narrow agent that can change a specified tile at each step, a turtle agent that can move to a next tile in the map grid and modify passed tiles, a wide agent that has control to edit any tile in the map, and a majority agent that suggests tile mutations if the majority of the other three agents have suggested that same mutation. The display of new suggestions is triggered by user interactions only. The user can accept any of these suggestions or override them by manually editing the grid for the current level. The tool also allows the user to fine-tune the performance of the AI models by providing two hyperparameters for the user to control. The step parameter defines the number of interactions for the manager of the models to recursively call itself, so that a higher step value allows for more than one modification to the map. The tool radius parameter controls how big is the window of tiles that the artificial agents can view and modify. By generating modifications one step at a time, this approach results in a creative process in which participants gradually build the level tile by tile.
Figure 4. The RL Brush tool (Delarosa et al., 2021). Top — the step and tool radius parameters; left — the toolbar to edit, erase, delete, play, and save the current level; middle — the editable grid of the current level and the four AI suggestions on top; right - the tileset. (Available at https://rlbrush.app/)
@@
Figure 5. The miWFC tool (Langendam and Bidarra, 2022). Left — a list with several example input images with the selected one displayed; middle - the panel for manipulating tile properties; right — the current image accompanied by a timeline indicating the progress of the generation and a number of controls for adding markers and navigating through the timeline. The screen for direct manual editing is accessible from the brush button in the bottom right corner. (Available at https://github.com/ThijmenL98/miWFC)
@@ 3.3.1. Pitako
Machado et al. (Machado et al., 2019a, b) developed a tool for designing tile-based games that applies the concept of a recommender system to suggest mechanics and dynamics learned from existing games. The system is built on Video Game Description Language (VGDL), which considers sprites as objects with a graphical representation and an associated behavior. The creative process can begin by adding sprites to the game in development. For instance, the user can define a sprite behavior as a ShootAvatar and then attach the image of the hero to it. Having a catalog of a vast number of games, the system can provide suggestions of sprites to the designer, who is free to accept them or not (Fig. 6). If the user accepts a recommended sprite, it is imported to the game sprite set along with any others it may carry (e.g., picking a sprite that can shoot will also import the sprite for the projectile behavior). In addition, the system also suggests interactions for pairs of sprites (e.g., enemy sprite kills avatar sprite), taking its cues from the existing games in the catalog. If the user selects a recommendation, the interaction is automatically implemented in the game.
This mixed-initiative approach primarily targets content creation within the class of game behaviors. However, it also extends to the game space and game design classes by enabling the placement of sprites on the level grid and the definition of game termination rules. Still, it only serves as a support tool for these two classes as the computational agent lacks initiative.
Figure 6. Sprite Recommendation in the Pitako tool (Machado et al., 2019a). Top — a menu featuring tabs for the several tasks within the game design process; left — the game’s sprite set, initially containing only a default background sprite; center — the sprite property editor, enabling the selection of an image and the configuration of its behavior; right — a list of the system-generated suggestions. (Available at https://github.com/tiago-lam/Recommender-System-for-Game-Design)
@@ 3.4.1. EvolvingBehavior
Partlan et al. (Partlan et al., 2022) introduced a tool for crafting non-player character (NPC) behavior by evolving behavior trees in Unreal Engine 4 (Fig. 7). The process begins with the human designer providing an initial draft of a tree and other materials, such as a library of pre-defined nodes and templates for generating additional nodes. The system can then modify and generate behavior trees, even if the initial tree is empty or incomplete. In order for the evolution of behavior trees to meet the goals of the human designer, they must define a fitness function by setting fitness measurement keys. The user also needs to define several parameters of the evolutionary process. Given that the evolved trees can be manually edited, in practice, the human designer can modify them and use them as input for a new round of evolution, engaging in an iterative process. However, this work was not selected for further analysis in section 4. This is because it is currently in a state where user interaction with the tool is too complex for a regular designer without coding skills and expert knowledge of the system to easily work with it. As acknowledged by the authors, it is challenging to define the fitness function and adjust all the parameters without prior knowledge. Additionally, the generation and subsequent simulation of behavior trees are time-consuming. The delay would likely discourage designers from participating in a new evolutionary process.
Figure 7. Some screens of the EvolvingBehavior tool (Partlan et al., 2022). Top left — an evolved tree; bottom left — template collection for starter behavior trees; bottom center — visualization of the NPCs’ behavior; top right — settings to define fitness functions; bottom right — mutation properties for the initial population. (Available at https://evolvingbehavior.npc.codes/)
@@ 3.5.1. Story Designer
The Evolutionary Dungeon Designer (EDD) is a mixed-initiative tool for designing 2D dungeons and the rooms that compose them. EDD is part of an ongoing research project and has been progressively expanded over several works (Baldwin et al., 2017a, b; Alvarez et al., 2018a, b, 2019, 2021). As a next step, Alvarez et al. (Alvarez et al., 2022) built an approach to design narrative structures integrated in EDD. Story Designer (Fig. 8) uses tropes in the form of nodes that can be interconnected in a graph to construct a complete narrative structure. The creative process begins with the system providing a basic sample graph HERO → CONFLICT → ENEMY. Users can manually edit the narrative graph by adding and connecting nodes representing characters, their roles, goals, and events. This triggers an evolutionary algorithm that generates alternative suggestions of narrative structure in real time that the user can inspect and choose to replace the current narrative design. The algorithm is an adaptation of MAP-Elites, which evaluates the set of solutions according to metrics such as interestingness and coherence. The system displays the suggestions in a matrix arranged by their values across specified dimensions of interest. Solutions with optimal fitness are shown in dark green, while dark red cells mean the best solution for those values is still not feasible. The human designer can visualize two dimensions at a time and switch between them as needed. As Story Designer is linked to the level design tool of EDD, the suggested narrative graphs are constrained by the existing content in the level’s map.
Figure 8. The Story Designer tool (Alvarez et al., 2022). Center — the editor for the main narrative graph; right — the suggestion grid, the list of dimensions that the user can choose for the grid, the suggestion they are hovering, and the selected suggestion; bottom — display of fitness metrics of the main graph and of the suggestions. (Available at https://github.com/mau-games/eddy)
@@
Figure 9. The 1001 Nights game (Sun et al., 2022). Left — a dialogue interface containing text blocks from the story told by the player and the King, along with a text box at the bottom for the player to write; right — a screen featuring the game characters, in which the player can collect the generated items and add them to the inventory at the bottom; bottom right corner: the button to initiate the battle. (Available at https://cheesetalk.itch.io/one-thousand-and-one-night)
@@ 3.6. Game Design
There are few MI-CCy approaches that address high-level concepts, such as the rules and objectives of a game. Most existing tools are instead tailored for a specific game or game genre. So far, when the creative task involves the ideation of a novel game, it is still uncommon to grant decision-making authority to the computational agent, relegating them to a passive role in the process. Still, there are some works that deviate from this pattern. Among the tools we have reviewed, CADI and Gamika fall into this category but were not selected. CADI (Mobramaein et al., 2018) is a tool that employs a conversational interface for agents to design variations of the game Pong. Gamika (Nelson et al., 2017) is a system for creating 2D physics-based games for mobile devices, which parametrizes the design space into 284 features that the user can manipulate.
@@ 3.6.1. Germinate
For instance, a user interested in creating a stress-themed game might define entities such as “Friend” and “Work”, a resource called “Stress”, and the relationships and triggers they want to showcase. This example is illustrated in Fig. 10. The user can explore the several generated games to assess if there are any interesting results and then make adjustments to the properties to better reflect their intent, allowing for iterative refinement of the game. This approach is focused on starting from high-level rhetorical goals and then working on details to meet those goals. To articulate the conceptual essence of the game, Germinate also relies on the creation of a game system and game behaviors.
Figure 10. The Germinate tool (Kreminski et al., 2020). Left — properties specified by the user; center — a playable version of one of the generated games (a single game is shown at a time); right — the properties of the generated game. (Available at https://github.com/ExpressiveIntelligence/Germinate)
@@ 4. Mixed-initiative Classification
Deterding et al. (Deterding et al., 2017) emphasize how the human and the computer iteratively respond to each other through a variety of possible actions. Following this idea, Spoto and Oleynik (Spoto and Oleynik, 2017) introduced a framework that allows the comparison of mixed-initiative systems regarding the flow of the creative process. The flow is visualized through a graph that considers seven subprocesses of the creative process, namely ideate, constrain, produce, suggest, select, assess, and adapt, which can be executed by both the human and the computer. The graph illustrates the sequence of actions through directional links, usually outlining internal interactive loops. Muller et al. (Muller et al., 2020) expand on this framework, addressing mixed-initiative scenarios with generative AI systems. Grabe et al. (Grabe et al., 2022) also adapt the framework to co-creativity with Generative Adversarial Networks (GANs), identifying four primary patterns of interaction between humans and GANs. Focused on interaction design in co-creative systems, Rezwana and Maher (Rezwana and Maher, 2021, 2023) presented COFI, a framework that describes several interaction components to represent the space of possibilities for interaction between humans and AI and with the shared creative content. We consider these frameworks to be valuable, making it possible to distinguish between approaches in terms of how the creative process admits the initiative of both agents involved and how they interact with each other. However, they make little to no mention of the degree of mixed-initiative, the real impact each agent has on the creative process, and how much their collaboration is facilitated.
@@ 4.1. MI-CCy Quantifier: A Framework for Mixed-Initiative Co-Creativity Quantification
Liapis et al. (Liapis et al., 2016) consider it ingenuous to expect human and computational agents to be equal partners in the creative process. We admit that participants involved in co-creation cannot be exactly equal throughout the various stages of the process, which is also true for human-human co-creativity, as each person has their own personality, knowledge, capabilities, and past experience. Nonetheless, we argue that such should be explored as one of the possible settings for future mixed-initiative tools. Liapis et al. (Liapis et al., 2016) reiterate that current approaches favor human control, with the user having the final say on the decisions taken. We hope that new approaches can challenge this view. As mentioned earlier, a high degree of human initiative combined with a low degree of computational initiative prevents the designer from fully leveraging the unique capabilities of AI systems. Conversely, a very high degree of computational initiative constrains the designer’s creativity and control. It is our belief that working as colleagues in a process in which collaboration is continuous and facilitated allows the human and the machine to learn from each other productively and can maximize the capabilities of both. This perspective stems from the observation that the most fruitful co-creative interactions often occur when both agents are able to take initiative, influence the direction of the creative outcome, and respond meaningfully to each other’s contributions. By allowing humans and AI to work as equals in the creative process, we can foster a richer and more dynamic collaboration that goes beyond the restrictive roles of mere assistance or automation. This enables more creative exploration and the emergence of surprising outcomes.
With this perspective in mind, we present the MI-CCy Quantifier, a framework for comparing systems with respect to the degree of mixed-initiative co-creativity. The classification is made independently of the process flow. In this way, the analyzed parameters are generalized and applicable to any system, even if the processes are very different in their form. This model allows for a more direct visualization of which agent has a greater impact and how seamless their collaboration is. Figure 11 shows the framework representing a scenario with a very high degree of mixed-initiative co-creativity. We emphasize that the MI-CCy Quantifier is not intended as a tool for rigorous classification, but rather as a means to support structured reflection on the mixed-initiative approach one seeks to develop.
@@
We introduce four parameters that allow us to classify the influence that the human and the computer have on the creative process: Initial Setting, Initiative, Evaluation, and Final Decision. Each of these parameters is assessed along a spectrum, with the human at one end and the computer at the other. The spectrum is divided into 13 levels in total, covering 7 levels between the human end and the middle (perfect balance) and 7 levels between the middle and the AI end. We consider that the more equal (towards the middle of the spectrum), the higher the degree of mixed-initiative. The framework has three more parameters that classify the approach on how much it promotes seamless collaboration (meaning that collaboration is facilitated, continuous, frictionless, and ideas are easily blended). They are Task Assignment, Intervention Pace, and Explainability. Each of these parameters comprises its own scale, which is displayed vertically. The scale is divided into 7 levels. We consider that the higher up the scale, the more seamless the collaboration between the human and the machine will be and, consequently, the greater the degree of co-creativity. These parameters are depicted in a visual scheme, where, in addition to the position on the scale of each criterion, we also establish a color chart for easier interpretation of the system’s MI-CCy level. Thus, if the evaluated parameter is green, it is because it aligns with the ideal conditions for strong MI-CCy, while intermediate colors leading to red denote varying degrees of compliance with the criteria.
@@
However, it should be noted that just because one system has a higher degree of mixed-initiative co-creativity over another does not mean that the approach is better than the other. It only tells us that the human agent and the computational agent are more equal partners and are equipped for a high level of collaboration. Such may or may not be desirable, depending on the intended applicability of the system. If the user already has an established idea and only wants help to execute it, they do not expect to have a system that wants to have a strong initiative in the conceptual phase of the process. Nevertheless, in this tutorial, we encourage researchers interested in developing new mixed-initiative co-creative systems to seek to implement an approach with a higher degree of MI-CCy since there is still little exploration in the current state of the art of the potential of the computer as a true colleague in the creative process. Only by providing the conditions for humans and artificial agents to collaborate in a seamless way can we properly assess how humans benefit from a more equitable partnership with AI systems.
We anticipate that a close and balanced collaboration between humans and AI can increase mutual adaptability, foster creativity and user engagement, and improve the integration of user goals and ideas with system-generated content. However, we must also be mindful of the challenges it poses. To work effectively as colleagues, we need to provide clear communication channels and an intuitive interface, as well as manage initiative and control thoughtfully. Poorly implemented approaches can overwhelm the user or lead to a reduced sense of agency. For this reason, establishing an equal relationship must be a deliberate decision that is appropriate to the creative context. Moreover, humans and AI have different capabilities that must be equated in order to provide a peer status with reciprocal influence that still accommodates different competencies.
It is worth clarifying that throughout this tutorial, we refer to both human and computational agents as capable of learning and exhibiting creativity in a practical sense. Drawing on Boden’s (Boden, 2004) taxonomy of creativity, we emphasize the importance of supporting transformational creativity in co-creative contexts, where either agent can contribute to altering the conceptual space or the creative direction. Our perspective aligns with the concept of symmetric alternating co-creativity (Kantosalo and Toivonen, 2016), in which both human and AI agents take turns in creating an artifact that satisfies both parties, and can use transformational creativity to handle conflicts and adapt to a new situation. While computational creativity and learning are fundamentally different from their human partners, we use these terms to describe behaviors that play comparable roles in MI-CCy — such as producing novel content, responding to feedback, or adapting to a partner — and do not imply cognitive equivalence between humans and machines. The MI-CCy Quantifier is thus meant to help reflect on how initiative is shared between agents and how easily they collaborate, regardless of the underlying nature of their reasoning.